. A certain wire, 3 m long, stretches by 1.2 mm when under tension of 200 N. By how much does
an equally thick wire 6 m long, made of the same material and under the same tension, stretch?
A. 4.8 mm B. 2.4 mm C. 1.2 mm D. 0.6 mm
1
Expert's answer
2020-05-12T09:56:56-0400

"Y=\frac{TL}{\Delta LA}"


As in both wire material is same hence the value of young's modulus will be same in both the cases

hence equating

"\frac{200\times3}{1.2\times10^{-3}\times A}=\frac{200\times6}{\Delta L\times A}"


"\Delta L=2.4\times10^{-3}m=2.4mm"